# Rethinking Inter-LoRA Orthogonality in Adapter Merging: Insights from Orthogonal Monte Carlo Dropout

### v1

@@ 4.3 Merging LoRAs from Community Sources
In addition to the controlled quantitative experiments presented in the previous section, Fig. 3 shows results from merging LoRAs downloaded from the community. Since our merging operates in the hidden space, LoRAs with different ranks can also be freely combined. These results demonstrate that our proposed method requires no additional fine-tuning and generalizes well, while further highlighting that orthogonality plays a limited role in LoRA merging.
@@ 5.1 Interference as a Source of Image Quality Gains
Direct Merge achieves substantially higher image quality than the other methods, indicating that the interactions arising from direct addition of LoRAs can, in fact, enhance visual quality. This observation calls into question the general assumption that orthogonality is always more advantageous for merging.
@@ 5.2 User Study: Isolating Orthogonality
While participants tended to favor higher-quality images, the Dropout vs Orthogonal case in Table 1 demonstrates that, under comparable image quality, enforcing strict orthogonality does not improve user-perceived concept fidelity.
@@ 5.3 Relationship with Zip-LoRA
We do not question the effectiveness of Zip-LoRA [7], but note that its performance likely stems from additional fine-tuning rather than orthogonality. Zip-LoRA does not guarantee orthogonality; it only encourages it through an inner-product penalty in the loss.
Direct Dropout Orthogonal $\uparrow$ MUSIQ [19] 62.8063 59.5810 59.6875 $\uparrow$ ARNIQA [20] 0.7529 0.7398 0.7423 $\uparrow$ NIMA [21] 4.2824 4.1120 4.1148 $\uparrow$ TRes [22] 87.3702 84.0270 84.1839 $\uparrow$ DBCNN [23] 0.6647 0.6419 0.6427
Table 2: Image Quality.
@@ 6 Conclusion
We introduced Orthogonal Monte Carlo Dropout to enforce strict orthogonality in LoRA merging. Our experiments show that orthogonality alone offers little benefit, as it neither yields semantic disentanglement nor improves compositionality, a negative result that we believe is valuable for guiding future research.

### v2

@@ 5.5 Merging LoRAs from Community Sources
In addition to the controlled quantitative experiments presented in the previous section, Figure 5 shows results from merging LoRAs downloaded from the community. Since our merging operates in the hidden space, LoRAs with different ranks can also be freely combined. These results demonstrate that our proposed method requires no additional fine-tuning and generalizes well, while further highlighting that orthogonality plays a limited role in LoRA merging.
Table 2: Image Quality.
Direct Dropout Orthogonal $\uparrow$ MUSIQ (Ke et al.,, 2021) 62.8063 59.5810 59.6875 $\uparrow$ ARNIQA (Agnolucci et al.,, 2024) 0.7529 0.7398 0.7423 $\uparrow$ NIMA (Talebi and Milanfar,, 2018) 4.2824 4.1120 4.1148 $\uparrow$ TRes (Golestaneh et al.,, 2022) 87.3702 84.0270 84.1839 $\uparrow$ DBCNN (Zhang et al.,, 2020) 0.6647 0.6419 0.6427
@@ 6 DISCUSSIONS
6.1 Merging LoRA weights or merging LoRA outputs?
Throughout this paper, we have referred to “merging LoRA modules.” In practice, however, our method operates by merging the outputs of LoRA modules. For clarity, note that directly merging the weights of LoRAs via a weighted sum is mathematically equivalent to merging their outputs, due to the linearity of matrix multiplication:
$(w_{1}\Delta W_{1}+w_{2}\Delta W_{2})h=w_{1}\Delta W_{1}h+w_{2}\Delta W_{2}h.$
6.2 Controlled Variables
In the three methods illustrated in Figure 1, one key controlled variable is the expected scale of the combined vectors. As shown in (1), although $z\sim\text{Ber}(1-p)$ reduces the expected scale by a factor of (1p)(1-p), dividing by (1p)(1-p) restores the expectation to 1. Consequently, as demonstrated in our redundancy experiments, even with dropout masking 90% of the LoRA activations, controlling the expected scale ensures that semantic preservation remains intact.
Another critical controlled variable is the initial noise in diffusion-based image generation. As noted by Guo et al., (2024) and Miao et al., (2025), the initial noise has a substantial impact on the final image. To the best of our knowledge, our study is the only one in this line of work that strictly controls the initial noise, which is essential for isolating and evaluating the true effect of orthogonality on semantic preservation.
@@ 6.3 Interference as a Source of Image Quality Gains
Direct Merge achieves substantially higher image quality than the other methods, indicating that the interactions arising from directly adding LoRAs can, in fact, enhance visual quality. This finding suggests that interference between LoRA modules is not always harmful; in some cases, it may introduce constructive interactions that improve perceptual fidelity. Such an outcome challenges the prevailing assumption that enforcing orthogonality is inherently more advantageous for merging. Instead, our results highlight a more nuanced trade-off: while orthogonality helps eliminate interference at the representational level, it may simultaneously suppress beneficial interactions that contribute to visual quality. This observation motivates a re-examination of the role of orthogonality in adapter merging, suggesting that effective strategies may need to balance interference reduction with the potential gains brought by controlled interactions.
@@ 6.4 User Study: Isolating Orthogonality
While participants generally preferred higher-quality images, the Dropout vs. Orthogonal comparison in Table 1 shows that when image quality is comparable, enforcing strict orthogonality provides no observable benefit in user-perceived concept fidelity.
@@ 6.5 Relationship with Zip-LoRA
We do not dispute the effectiveness of Zip-LoRA (Shah et al.,, 2023); however, we note that its performance is more likely attributable to the additional fine-tuning it requires rather than to orthogonality itself. Importantly, Zip-LoRA does not enforce strict orthogonality. Instead, it merely encourages orthogonality by introducing an inner-product penalty into the loss function, which provides no theoretical guarantee of orthogonal representations.
6.6 What Do People Expect from Orthogonality?
Orthogonality is an elegant mathematical property, but its semantic implications are often vaguely defined. For example, should orthogonality lead to better disentanglement of semantics, or to better fusion of semantics? Although these notions lack precise mathematical definitions, disentanglement and fusion are often seen as contradictory. Thus, claims such as “orthogonality improves disentanglement” or “orthogonality improves fusion” are largely intuitive rather than rigorously grounded. Our findings demonstrate that, in the context of LoRA merging, orthogonality provides no tangible benefit. This suggests that the community should adopt a more cautious perspective when attributing semantic advantages to the mathematical property of orthogonality.
@@ 7 CONCLUSION
We introduced Orthogonal Monte Carlo Dropout to enforce strict orthogonality in LoRA merging without extra computational cost. While our theoretical analysis confirms runtime orthogonality and our experiments show the redundancy of LoRAs across both controlled and community settings, the results reveal a clear negative finding: orthogonality alone brings little benefit, as it neither improves semantic disentanglement nor compositionality, and in some cases even reduces image quality compared to direct merging. This suggests that the field should rethink the assumed semantic advantages of orthogonality in adapter merging and explore alternative strategies that balance interference reduction with the preservation of constructive interactions.